Write the slope-intercept equation for the line that passes through (2, -2) and (-1, 10
Answer by mananth(16949)

x1 y1 x2 y2
2 -2 -1 10

slope m = (y2-y1)/(x2-x1)
(10- -2)/(-1 -2)
( 12 / -3 )
m= -4

Plug value of the slope and point ( 2 , -2 ) in
Y = m x + b
-2.00 = -8 + b
b= -2 - 8
b= -10
So the equation will be
Y = -4 x -10
m.ananth@hotmail.ca